# Design and deployment of a fast neural network for measuring the properties of muons originating from displaced vertices in the CMS Endcap Muon Track Finder

### v2

@@ 2 Design and Hardware Implementation
Therefore an NN-based solution is chosen to be implemented in Run 3 EMTF system that targets measurement of and of displaced muons. Two separate NNs are developed in software to target and separately. Each of the NN models have three dense layers, and equally split half of the total nodes of the complete model as seen in Fig. 1. The two models are then stitched together to form a single NN model before being implemented into the FPGA. The models are trained on simulations of LLPs that decay to muons made with standard CMS software emulators.
@@ 3 Deployment for Operations and Expected Performance
The displaced NN was deployed for CMS operations on 26th June 2023, making it the first NN running online in CMS L1T FPGAs during data taking. The commissioning of the NN was performed during the 2023 data taking year using monitoring triggers such as the one seen in Fig. 3, and the first set of triggers making use of the NN based and assignment were deployed for data taking in 2024. Fig. 3 also shows the expected performance of the displaced NN and prompt BDT algorithms on a simulated sample of displaced muons produced with an earlier design of the NN model from 2022 [5]bbbA more recent version of this plot was published in CMS-PAS-EXO-23-016.

### v3

@@ 2 Design and Hardware Implementation
Therefore an NN-based solution is chosen to be implemented in Run 3 EMTF system that targets measurement of and of displaced muons. Two separate NNs are developed in software to target and separately. Each of the NN models have three dense layers, and equally split half of the total nodes of the complete model as seen in Fig. 1. The two models are then stitched together to form a single NN model before being implemented into the FPGA. The models are trained on simulations of LLPs that decay to muons made with standard CMS software emulators. The trainings are done separately using M events, each containing 4 muons, both minimizing similar log-cosh loss functions over 300 epochs. The 29 input features are calculated for each muon track, and include of the track measured from the beam axis, differences of and angles and their signs between each stub used to build the track, and a parameter that measures the bending recorded by each stub.
@@ 3 Deployment for Operations and Expected Performance
The displaced NN was deployed for CMS operations on 26th June 2023, making it the first NN running online in CMS L1T FPGAs during data taking. The commissioning of the NN was performed during the 2023 data taking year using monitoring triggers such as the one seen in Fig. 3, which shows that the deployed NN works with a rate vs pile-up (PU) profile as expected from simulations. Fig. 3 also shows the expected performance of the displaced NN (solid stars) and prompt BDT algorithms (hollow squares) on a simulated sample of displaced muons produced with an earlier design of the NN model from 2022 [5]bbbA more recent version of this plot was published in CMS-PAS-EXO-23-016. The different colors show different regions: 1.2 < || < 1.6 (black), 1.6 < || < 2.1 (red), and 2.1 < || < 2.5 (blue). The first set of triggers making use of the improvements from the NN based and assignment were deployed for data taking in 2024. These new triggers increase the CMS trigger acceptance to highly displaced muons ( cm), which substantially improve signal sensitivities to models with LLPs that decay to muons in the final state.